OpenStudy (anonymous):

find the inverse of the function f(x)=x^3-8 (urgently needed plz)

OpenStudy (saifoo.khan):

y = x^3 - 8 Make x the subject.

OpenStudy (anonymous):

i have four options to choose from Not a one-to-one function 2) f^-1(x)= 3sq x+8 3) f^-1(x)=3sq x-8 4) f^-1 (x) = 3 sqx=8

OpenStudy (anonymous):

which one would be the right option

OpenStudy (anonymous):

take f^-1(y)=x y=x^3-8 \[\sqrt[3]{y+8}=x\] so f^-1(y)=\[\sqrt[3]{y+8}\]

OpenStudy (anonymous):

2) f^-1(x)= 3sq x+8 ur answer

OpenStudy (anonymous):

thankyou so much.

OpenStudy (anonymous):

medal plz :)

OpenStudy (anonymous):

im new to this thing. how do u do that?

OpenStudy (anonymous):

i'm new too. You give a medal by clicking on "best response".

OpenStudy (anonymous):

yep i got it thankyouuu! :D

OpenStudy (anonymous):

But I don't think I can give myself a medal...

OpenStudy (anonymous):

yep i think so too

OpenStudy (anonymous):

a medal for you from me then...

OpenStudy (anonymous):

awww thankyou